an angles measure is 6 degrees more than 3 times the measure of its complement. find the measure of the angles
Answer by cleomenius(959)

x + 6 + 3x = 90
4x = 84
x = 21 degrees, = The smaller angle.
3(21) + 6 = 69 degrees, the larger angle.
21 + 69 = 90 degrees, this does check.
Cleomenius.